OpenStudy (anonymous):

how does -(2x+y)/(x-2y) = (2x+y)/(2y-x)

OpenStudy (anonymous):

what do you have to find ?

OpenStudy (anonymous):

-(a-b) is same as -a+b which is same as (b-a)

OpenStudy (anonymous):

hahahaahah... NOW I GET IT ! \[\LARGE -\frac{2+xy}{x-2y}\;\;\; \text{can be written as...} \frac{2+xy}{-(x-2y)}\] and now you have SUROJ answer!

OpenStudy (anonymous):

-(2x+y)(2y-x)=(2x+y)(x-2y) \[-(4xy+2y ^{2}-xy-2x ^{2})=2x ^{2}-4xy+xy-2y ^{2}\] 0=0 :D:D:D:D:D

OpenStudy (callisto):